The probability distribution for the number of cars sold during a car dealer is given below:
X
0
1
2
3
4
5
(X)
0.18
0.39
0.24
0.14
0.04
0.01

The probability that at least 3 cars were sold on a particular day is:

Answer by ikleyn(52873)

.

    P(at least 3) = P(3) + P(4) + P(5) = 0.14 + 0.04 + 0.01 = 0.19.    ANSWER

Solved and explained. And completed.